OpenStudy (lovelycharm):

In December 1999, the exchange rate of Mexican pesos to American dollars was 9.52 to 1. You paid 450 pesos for a jacket. Use the following verbal model to find the price of the jacket in dollars.

OpenStudy (lovelycharm):



OpenStudy (destinymasha):

9.52/1 = 450/x 9.52x = 450 , x = $47.27

OpenStudy (beccaboo333):

Hi, next time please ask in the correct section. (mathematics)